### v1

# KIX: A Metacognitive Generalization Framework

Abstract
Humans and other animals aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high level knowledge acquired over time. But artificial agents are more of a specialist, lacking such generalist behaviors. Artificial agents will require understanding and exploiting critical structured knowledge representations. We present a metacognitive generalization framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects leveraging type space facilitate the learning of transferable interaction concepts and generalization. It is a natural way of integrating knowledge into reinforcement learning and promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
1 Introduction
Figure 1: Learning for generalist behaviors necessitates building a natural yet abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach captures the inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
Flexibility, ability to actively think and evaluate options –learning and reasoning to enable adaptation– are hallmarks of human intelligence. Core to this intelligence is how easily people can find natural decomposition in world environments and tasks, enabling flexible, generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge-centric metacognitive approach.
Meta-reasoning about next actions, action values, and possible next states facilitates flexible and transferable decision making for novel situations, problem solving, concept learning, and knowledge acquisition. This reasoning involves more than just memory lookup. Meta-reasoning is more broadly sense-making in a knowledge system – enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in different ways, and ground their interactions for both short and long-term goals. Metacognitive reasoning appears to be central to generalization. In contrast, meta-reasoning systems in artificial intelligence are more focused on resource allocation problems. Although resource allocation underlies efficient functioning, we need to ideate metacognitive generalization, figure 1, to enable generalist behaviors in artificial agents. Generalist behaviors have key characteristics which both constitute a basis of devising a metacognitive framework and compartmentalize the behaviors in a useful way.
Our proposal builds on three main principles for solving complex tasks in a flexible manner. Natural decomposition and abstraction: World environments are naturally composed of entities and categorized by their types. It is similar to how a natural hierarchy of objects and classes is formed. Agents interact with entities in the instance space, which inherit properties and relationships from a higher-level type space. Interactions: World environments are essentially knowledge graphs that represent the underlying objects and their relationships. In order to complete its tasks and achieve its goals, an agent’s metacognitive process employs these knowledge graphs to create a plan that specifies which entities it must interact with and what interactions it should have. Composition of interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at the type-space and instance levels as needed for quick adaptation. Integrating these principles into metacognitive learning, reasoning, and control provides levers and connecting points for instilling generalist behaviors in agents.
Our proposed meta-cognitive framework captures the essence of generalist behaviors and may have far-reaching implications across domains. In this article, we provide a description of the key components of the proposed framework, which exemplify generalist characteristics followed by experimental evaluation, as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
2 Background
In this section, we briefly describe closely related work as well as the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
Related work Compositionality and abstractions are central for generalization learning. Abstraction requires a method for grounding representations to the actual objects. In previous work, Veerapaneni et al. (2020) presented an object-centric model in which object representations were constructed from raw observations and local simpler functions with uniformity, and then used for prediction and planning. Similarly, Vijay et al. (2019) used a symbolic object-centric representation to study the effect of adding prior knowledge through relations in an object graph. These works only consider object space, which inhibits the abstraction of objects and relational structures that occur naturally in the real world and are used for making decisions. Some of the main questions are what an abstraction should entail and how it can be used to compose interactions that can be used as building blocks for leveraging natural relational structures. We attempt to address these questions, which have been largely overlooked in previous works. Moreover, what does natural abstraction mean? It implies preserving relational higher-order structures and abstractions that occur naturally in the real world. Our proposal preserves abstract knowledge in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is accomplished by learning over knowledge graphs, with neighborhoods aggregated until stability is reached. Different from previous works, that’s what our proposal entails.
Other methods emphasize aspects which are different from ours, such as encouraging exploration through intrinsic reward or information gain Ten et al. (2022); Oudeyer et al. (2007); Pathak et al. (2017), or motivation Moerland et al. (2023); Zhu et al. (2023); exploring for a variety of experiences by developing internal goals for empowerment Klyubin et al. (2005) ; using a curriculum to progress from easier to more difficult tasks Soviany et al. (2022); Campero et al. (2020) or inducing a learning hierarchy either based on teacher signals Portelas et al. (2020) Pateria et al. (2021); Hutsebaut-Buysse et al. (2022) or rewards Furelos-Blanco et al. (2023). These methods focus on the execution space, influencing reward signals for learning policies in one way or another but largely ignore the higher order structures and relations that objects in the environment have. Our work distinguishes itself by incorporating natural relational structures in the form of knowledge graphs and allowing such meta-level knowledge structures to guide interactions that govern low-level execution.
Meta-reasoning Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, meta-cognitive reasoning in its current form as a resource manager is insufficient to exhibit characteristics of generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
Graph neural networks A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes. Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$. One typical task is to predict the label or value of the entire graph. Graph neural networks (GNN) (Gilmer et al. (2017); Hamilton et al. (2017) and others) use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$. Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
$h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable, permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which appears to improve static attention problems in graph attention layers.
Policy gradient methods We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992) , directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor critic methods using clipped objectives. We adapt PPO for learning policies.
3 Model
Our proposal for meta-cognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects. Meta-policy recommends localized agent-object interactions. The agent executes the interactions following interaction policies and constructs a compositional solution. Here, we discuss the main ideas of the framework (see figure 2) and conjecture that it aptly exhibits task generalization characteristics.
Figure 2: The meta-cognitive generalization framework consists of three levels: meta-level knowledge (K), interaction (I), and execution (X), which are based on levels of abstraction and computation to provide generalization capabilities. The knowledge or type space contains type space knowledge graphs representing higher order abstractions. The instance space contains object-level instance graphs and functions as an intermediary. The execution space is the lowest level abstraction in which the agent acts on meta-level recommendations. Computations proceed through recognition, recommendation, execution, and interactions to gather experiences for updates to interaction and meta policies.
Knowledge An agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ in the environment $\Xi$. The environment consists of objects, which the agent’s perception systems associate with the entity types. Agent maintains its perceived knowledge in the form of knowledge graphs: a graph of objects and their relationships at the instance level, and a graph of entity types and their abstract relationships at the type level. A type space knowledge graph is an ordered pair $G_{K}=(E,\Lambda)$ that consists of a set of entity type nodes, $E=\{E_{1},E_{2},...,E_{m}\}$, and a set of edges or type level relations, $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. The same type of relation between objects and abstract entity types has the same meaning; for example, adjacency relations indicate that two objects are adjacent to each other, whether in instance or type space. Using the knowledge graph and meta-level action of the meta-policy, $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$, the meta-level of the agent determines how it will interact with an underlying object. To interact with an object, the agent activates an interaction type, $\lambda$, following the meta-policy guidance. For an interaction, the agent will perform a series of low-level actions. For example, if there is a door, the agent can activate ‘open the door’ at the abstract level, even though it requires many low-level actions such as moving and toggling. The agent uses an activation relation to keep track of which objects it wants to interact with and what type of interaction. Thus, the agent’s knowledge space $\mathcal{K}=(A,E,\Lambda,\lambda,G_{K})$ comprises of a type space knowledge graph, a set of entity types, abstract relations, and interactions.
Interaction The environment consists of objects, $e=\{e_{1},e_{2},...,e_{n}\}$, their types, $\text{type}(e_{i})\in E,\text{for}\ e_{i}\in e$, and is partially observable. Agent keeps track of the visible objects in its instance space and builds an instance graph, which is an ordered pair of nodes (observed objects) and edges (observed object relations), $G_{I}=(e,\Lambda_{I})$. Thus, the agent’s instance space, $\mathcal{I}=(A,e,\Lambda_{I},\lambda,G_{I})$, is made up of its instance space knowledge graph, relationships between objects, and available interactions. After constructing the knowledge graphs, the agent’s goal is to identify the objects for interaction which is done by the meta-level. Meta-level recommends an interaction goal for the agent. For the recommended interaction goals, the agent learns interaction policies for interacting with objects, as well as the relationships between interaction types and entity types. With the ability to interact with objects, the agent accumulates interaction knowledge to update the meta-level policy. Therefore, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals.
eXecution By pursuing interaction goals as directed by the meta-level, the agent learns to interact with objects and act at the execution level. The execution space for the agent is its partially observed view $X$. During an interaction, the agent takes actions $u$ on low-level states $s$. After learning both the interaction and meta policies, the agent acts in the environment, aided by interaction with objects. Therefore, unlike the models that are tightly coupled with low-level representations through end-to-end policies, agent actions are aided by meta-level knowledge-guided object interactions, providing a mechanism for generalist behaviors.
Mapping between execution space and instance space Assume that the agent has an object recognition system that can access its partially observed view $X$. The agent generates an instance-level graph using the observed objects. The function $\Phi:X\to G_{I}$ maps the agent’s object recognition to its instance graph by capturing observed objects and their relationships. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view. The mapping function then generates the corresponding instance graph.
Mapping between instance space and type space The function $\Psi:G_{I}\to G_{K}$ provides a mapping between instance and type graphs. If the agent encounters a novel object for which it does not have a corresponding typed entity, it must either create a new type in its type space or assign the object to a type that is closely related. In the framework, $\Psi$ is learnable and represents a belief of mapping objects to their types. However, in this article, we treat the mapping as a fixed mapping from object nodes in $G_{I}$ to their corresponding type nodes in $G_{K}$.
Proposal for activation object and interaction Using the meta-policy, the agent’s recommendation function $\Upsilon:G_{K},G_{I},\pi\to e_{i},\lambda_{p}$ recommends a candidate object $e_{i}$ for interaction and interaction goal $\lambda_{p}$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$, then activates the object by adding an activation relationship between the agent and the object to the instance graph, as well as a corresponding activation relation between the agent and the object’s type node in the type space graph $G_{K}$. Activated $G_{K}$ represents the activated candidate object and serves as the input state for meta-policy, which returns both the state value and meta-action probabilities. The state value represents the value of activating the respective object. Thus, the recommender proposes an object for interaction $e_{i}$ by sampling from the values of activating objects, and the proposed interaction $\lambda_{p}$ is sampled from the meta-action probabilities of the proposed object. Acting on these recommendations, the agent collects interaction experiences $G_{K},\lambda$, that is, interaction trajectories $G_{K}^{0}\curvearrowright G_{K}^{1}...\curvearrowright G_{K}^{t}$, and learns meta-level policy.
Interactions At each step, the agent activates the object $e_{i}$ for interaction $\lambda_{p}$, which is its interaction goal, $g^{i}_{p}=(e_{i},\lambda_{p})$. Following this interaction goal, the agent interacts with the environment. After each round of interaction, the status of the interaction is passed to the meta level. Using the meta-policy, the meta-level recommender suggests next target object and interaction. The process continues until the task is completed or the environment is terminated. Throughout the process, the agent collects experience trajectories at both the interaction and meta levels. The policies are then updated based on the collected experiences.
Meta-policy learning Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge type encoding for different types of relationships. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,
$h_{v}^{\prime}=\alpha_{v,v}\theta h_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{u,v}\theta h_{u}$
where $h_{v}^{\prime}$ is the updated embedded representation of node $v$, $\mathcal{N}(v)$ is the set of neighbors of node $v$, and $h_{u}$ is the representation of the neighboring node. $\theta$ are learnable weights shared by source and target nodes in the neighborhood, and the attention coefficients $\alpha_{u,v}$ are computed using weighted node features.
The hidden representation after the final graph attention layer is pooled for graph rollout, which is then passed to the fully connected actor and critic layers of the model. The meta policy $\pi_{\theta}(.|G_{K})$, parameterized by $\theta$, returns a state value and an interaction or meta-action, $\lambda$, distribution. The recommender selects a target object and an interaction $\lambda_{p}$, and the agent works through low-level interactions based on guidance. At the end of the interaction, the status of the interaction is returned to the meta level, and the agent collects the trajectory of meta-level experiences, $\mathcal{D}_{\tau_{m}}$. An estimate of advantages is computed as $A_{m}=r_{m}+\gamma_{m}V_{\pi_{\theta}}(G_{K}^{\prime})-V_{\pi_{\theta}}(G_{K})$.
Meta level PPO optimizes a surrogate loss:
$\mathcal{L}_{m}^{\text{actor}}(\theta)=-\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\min\left(r_{\theta}A_{m},\text{clip}(r_{\theta},1-\epsilon,1+\epsilon)A_{m}\right)\right]$
with $r_{\theta}=\frac{\pi_{\theta}(.|G_{K})}{\pi_{\theta}^{r}(.|G_{K})}$ being the ratio of probabilities of actions at meta level roll-out. The critic loss is the mean squared error between the estimated values and the returns:
$\mathcal{L}_{m}^{\text{critic}}(\theta)=\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\frac{1}{2}(V_{\pi_{\theta}}(G_{K})-R_{m})^{2}\right]$
Considering the mean entropy $\mathcal{H}_{\theta}$, the overall minimization objective at the meta-level becomes:
$\mathcal{L}_{m}(\theta)=\mathcal{L}_{m}^{\text{actor}}(\theta)-\zeta_{m}^{h}\mathcal{H}_{\theta}+\zeta_{m}^{c}\mathcal{L}_{m}^{\text{critic}}(\theta)$
where $\zeta_{m}^{h}$, $\zeta_{m}^{c}$ are the entropy and value coefficients, respectively. The model parameters $\theta$ are optimized as $\theta\leftarrow\theta+\alpha_{m}\nabla_{\theta}L_{m}(\theta)$, and the updated meta-policy is used for the next episode of learning.
Interaction policy learning Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately. When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$. The agent observes a low-level state $s$ from a convolutional neural network embedding, samples an action $u\sim\mu_{\phi_{p}}(.|s)$ using the respective interaction policy, acts in the environment, and collects the trajectory of experiences $\mathcal{D}_{\tau_{p}}$. The experiences collected are used to learn interaction policies. An estimate of advantages is computed as $A_{p}=r_{p}+\gamma_{p}V_{\mu_{\phi_{p}}}(s^{\prime})-V_{\mu_{\phi_{p}}}(s)$.
Each interaction level PPO optimizes a surrogate loss:
$\mathcal{L}_{p}^{\text{actor}}(\phi_{p})=-\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\min\left(r_{\phi_{p}}A_{p},\text{clip}(r_{\phi_{p}},1-\epsilon,1+\epsilon)A_{p}\right)\right]$
with $r_{\phi_{p}}=\frac{\mu_{\phi_{p}}(.|s)}{\mu_{\phi_{p}}^{r}(.|s)}$ is the ratio of probabilities of actions of policy and roll-out policy. The critic loss is the mean squared error between the estimated values and the returns:
$\mathcal{L}_{p}^{\text{critic}}(\phi_{p})=\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\frac{1}{2}(V_{\mu_{\phi_{p}}}(s)-R_{p})^{2}\right]$
Considering the mean entropy $\mathcal{H}_{\phi_{p}}$, the overall minimization objective becomes:
$\mathcal{L}_{p}(\phi_{p})=\mathcal{L}_{p}^{\text{actor}}(\phi_{p})-\zeta_{p}^{h}\mathcal{H}_{\phi_{p}}+\zeta_{p}^{c}\mathcal{L}_{p}^{\text{critic}}(\phi_{p})$
for each interaction $p$, where $\zeta_{p}^{h}$, $\zeta_{p}^{c}$ are the entropy and value coefficients, respectively. For each interaction policy, the model parameters $\phi_{p}$ are optimized as $\phi_{p}\leftarrow\phi_{p}+\alpha\nabla_{\phi_{p}}L_{p}(\phi_{p})$, and the updated interaction policies are used for the next episode of learning.
As a result, the agent follows a trajectory of activated interactions with objects and aids the formation of a task solution through meta-guided interactions. The proposed framework is generic to be applied to a wide range of domains.
4 Experiments
Experimental set up To assess the capabilities of the proposed agents, we require a challenging environment with ingredients for higher order reasoning to solve tasks. Our goal here is to investigate whether the introduction of meta-knowledge guided interactions benefits agents. MiniGrid world environments Chevalier-Boisvert et al. (2018) are suitable for our purpose. These are procedurally generated environments in which the agent interacts with objects such as doors, keys, balls, and boxes.
Figure 3: MiniGrid ObstructedMaze Full environment is an appropriate testbed because it is partially observable, requires complex reasoning, and new tasks can be created.
The environments are challenging due to partially observability, sparsity of rewards and random goal locations. The MiniGrid-ObstructedMaze-Full environment, figure 3, in particular, is a difficult environment with only a few claims for solving it. Due to the large state space size, stochasticity and partial observability, agents cannot memorize but must learn.
In the environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling. The original task is referred as Task 0. Furthermore, we identify the following additional task scenarios that were not seen during training but were presented to the agent during testing:
  • No context in training (Task 1) - a relation that the agent did not come across during training. It is simulated by placing the goal inside an empty box. During training, the blue ball was never in a box, although the agent encountered boxes and balls during training, but boxes contained keys.
  • Different context in testing (Task 2) - using a relation that the agent encountered during training but in a new context during testing. It is simulated by locking doors in the central room and placing keys directly in the room. During training, doors in the central room were never locked, nor was there direct access to keys.
  • Dynamic changes in the environment (Task 3) - when some event occurs, the environment changes. During training, the goal location remained unchanged. To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms.
Agent training
Initialize model weights $\theta,\phi$ , parameters, $T$ for  $t=1$ to $T$ do
       // collect exp for meta exp size K
       for $k=1$ to $K$ do
             make instance graph $G_{I}\leftarrow\Phi(X)$ map $G_{I}$ to $G_{K}$, $G_{K}\leftarrow\Psi(G_{I})$ proposal $(e_{i},\lambda_{p})\leftarrow\Upsilon(G_{K},G_{I},\pi)$ // 1st variant
             interaction exp following $\mu_{\lambda_{p}}$ // 2nd variant
             reaching exp following $\mu_{\text{reach}}$ if success then
                   interaction exp following $\mu_{\lambda_{p}}$
             end if
            collect exp $\mathcal{D}_{\tau_{p}}$ and compute $A_{p}$
       end for
      collect meta exp $\mathcal{D}_{\tau_{m}}$ and compute $A_{m}$ // Update policies, both variants
       update $\pi$ by optimizing $\mathcal{L}_{m}(\theta)$ for  $\lambda_{p}$ do
             update $\mu_{\lambda_{p}}$ by optimizing $\mathcal{L}_{p}(\phi_{p})$
       end for
      // 2nd variant only
       update $\mu_{\text{reach}}$ by optimizing $\mathcal{L}_{\text{reach}}(\phi_{\text{reach}})$
end for
Algorithm 1 KIX
We devise two variants, KIX1 and KIX2. Both variants learn meta-policy and interaction policies. However the second variant additionally learns a reachability policy, so it interacts with an object once it has reached it. All meta, interactions, and reachability policies are implemented as actor-critic PPO policies. Algorithm 1 provides a high-level overview of the implementation. The input state embeddings for interaction and reachability actor-critic networks are constructed by a convolutional network with three Conv2D layers with 16, 32, and 64 output channels, a max pooling layer added after the first convolution layer, and exponential linear units used as activation functions. The meta-level recommends various objects for interaction, so an activation indicator for the object of interest is added to a new input channel, creating a four-channel input. Actors and critics are fully connected networks, each with one hidden layer. Meta-level policies are learned over type space knowledge graphs. In a type space graph, node object type information is encoded as an attribute vector. Edges in type space graph represent type-level relationships between nodes, such as "visible," "adjacent," "carrying," and "activated." A one-hot encoded edge attribute vector is used to encode these relationships. Embeddings for meta-level actor-critic are provided by two graph attention convolution layers, each with 16-dimensional embeddings and four attention heads. During message passing, these layers share weights between source and target. The final graph attention convolution layer is used for graph readout with a global add pool. Meta-level actors and critics are fully connected networks, with the critic learning a single value and the actor learning the probability of taking a meta-level action, such as ’pickup’, ’drop’, ’reveal’, ’open’, or ’open with key’. The meta-level actions serve as interaction goals. The environment rewards the agent with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps. For interaction policy learning, the agent receives a reward of 1 when it has a successful interaction. At the meta-level it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal. A base agent is also implemented, with a similar network and a PPO policy, and it receives environmental rewards. Grid search is used to select hyperparameters. All the agents are trained for the same number of environment steps using 32 parallel workers.
Figure 4: The return profiles of agents performing various evaluation tasks show that the proposed agents have higher returns on all tasks and benefit from meta-guided interactions.
Meta-guided interactions aid agents Our proposal is motivated by the idea that, rather than acting directly at the execution level, meta-level guided interactions with objects aid agents in generalizing while performing tasks. Because the agent’s environmental return is determined by the number of steps taken to complete the task, completing tasks in fewer steps yields a higher return. Therefore, the distribution of environmental returns is a reliable indicator of whether meta-level interaction recommendations benefit agents. We evaluate the distribution of environmental returns for each agent across all four tasks. For each task, the trained agents are run for 12K episodes by parallel workers, and their environmental rewards are recorded. In order to understand the effects of meta-guided interactions and to ensure a fair comparison between agents where differences in success rates are not a factor, we examine the top $k$ returns with $k=100$. Figure 4 illustrates the return profiles of agents. It is clear from the distribution of returns that the proposed agents have received a higher return on all tasks than the base agent, furthermore, this is even more so for dynamic tasks like task 3. It demonstrates that meta-guided interactions not only help agents complete their task goals in fewer steps, but also aid them to perform a wider range of tasks in static or dynamic environments. Furthermore, the higher returns of the second variant demonstrate that agents benefit from the introduction of reachability as a separate interaction. In the absence of a separate reachability policy, reaching is implicit in each interaction policy; however, separating allows for a shared learning of reachability followed by interactions which could be beneficial particularly in dynamically changing environments.
Figure 5: Wasserstein distances between traversal distributions, when agents complete tasks 1-3, and the trained task 0, provide comparative insights.
Agents profile While performing the tasks, agents interact with objects in the environment according to the recommended interaction policies; as a result, agents navigate and visit rooms. Because interactions are recommended by the meta-level, which is only trained on the training task, tracking agents’ room visits across all tasks provides an insight about their learning. Let $p_{0}$ be the distribution of agent room visits for trained task 0, and $p_{i},i\in\{1,2,3\}$ be the distributions for tasks 1–3. For both variants of the agent, we then compute the Wasserstein distance, a metric from optimal transport theory, between the visit distribution for tasks 1-3 and the trained task 0, as shown in figure 5. The distribution of visits to each agent is expected to be different from the training task, due to changes in the evaluation tasks. However, our interest is in how differently the two variants behave in evaluation tasks, as indicated by the computed metric. In task 1, the goal room remained unchanged, but a new relation was added to the evaluation. For both variants, the computed metric value for task 1 is similar, indicating that they transfer their learning to this task in a similar manner. However, there is a noticeable difference for tasks 2 and 3. In task 2, doors in the central room are locked with keys kept in the room - a static task, whereas in task 3, the goal is dynamically relocated based on observance - a dynamic task. The computed metric for the dynamic task for the second variant is smaller than the first variant, implying that its evaluation visit distribution is closer to that of the training task. It suggests that the learning in the second variant can better capture dynamic changes. When combined with the return profiles, it reinforces that having an additional reachability policy benefits the agent in a wider range of tasks, particularly dynamic tasks, and improves generalization.
5 Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist and flexible behaviors necessitate abstraction and composition, but current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, natural abstractions and compositions are represented and grounded using type space knowledge graphs which also preserve relational knowledge between objects. Second, rather than directly acting in the raw environment, interactions with objects play an important role in how an agent navigates the environment to complete tasks. To accomplish this, we propose an intermediary interaction level. The meta-level of the agent learns from type space knowledge graphs and recommends what objects to interact with, as well as what interaction should be. In this sense, our model also functions as an interaction goals generating machine. Third, interaction policies learned through interaction goals serve as a repository of policies for the agent to use in the execution space, providing for the necessary loose coupling for generalization via interactions as intermediaries. The agent’s meta-level recommends interactions based on knowledge structures and learned meta-policy, and the execution level applies interaction policies to complete tasks. We demonstrated that metacognitive knowledge guided interactions aid agents in generalization. Our framework is modular and adaptable, with the potential to have implications across domains through the induction or integration of knowledge structures, whether extrinsic, intrinsic, or derived, as well as control variables at the meta or interaction level.