### v2

# KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization

Abstract
People aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time. But artificial agents are more like specialists, lacking such generalist behaviors. Artificial agents will require understanding and exploiting critical structured knowledge representations. We present a metacognitive generalization framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects leveraging type space facilitate the learning of transferable interaction concepts and generalization. It is a natural way of integrating knowledge into reinforcement learning and is promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
1 Introduction
Figure 1: Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach exploits inherent abstractions and relationships between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
Flexibility, ability to actively think and evaluate options –learning and reasoning to enable adaptation– are hallmarks of human intelligence. Core to this intelligence is how easily people can find natural decomposition in world environments and tasks; this enables flexible, generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge and interaction-centric metacognitive approach.
Meta-reasoning about next actions, their values, and possible next states facilitates flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition. This reasoning extends beyond mere memory lookup. Meta-reasoning is more broadly sense-making in a knowledge system - enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in diverse ways, and ground their interactions for both short- and long-term goals. Metacognitive reasoning appears to be central to generalization. In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems. While resource allocation underpins efficient functioning, it is crucial to ideate metacognitive generalization, Figure 1, to enable generalist behaviors in artificial agents.
Our proposal builds on three main principles for solving complex tasks flexibly. Natural Decomposition and Abstraction: World environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes. Agents interact with entities in the instance space, which inherit properties and relationships from a higher-level type space. Interactions: World environments are essentially knowledge graphs representing underlying objects and their relationships. An agent’s metacognitive process employs these knowledge graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals. Composition of Interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at both the type level and the instance level, enabling quick adaptation. Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents.
The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains. In this article, we describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
2 Background
In this section, we briefly describe closely related work and the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
Related work
Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects. In previous work, Veerapaneni et al. (2020) presented an object-centric model where object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning. Similarly, Vijay et al. (2019) used a symbolic object-centric representation to study the impact of adding prior knowledge through relations in an object graph. These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making. Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks leveraging natural relational structures. We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is achieved through learning over knowledge graphs, with neighborhood aggregation until stability is reached. This distinguishes our approach from previous works.
Other methods emphasize different aspects, such as encouraging exploration through intrinsic reward or information gain Ten et al. (2022); Oudeyer et al. (2007); Pathak et al. (2017), motivation Moerland et al. (2023); Zhu et al. (2023), developing internal goals for empowerment Klyubin et al. (2005), using a curriculum to progress from easier to more difficult tasks Soviany et al. (2022); Campero et al. (2020), or inducing a learning hierarchy based on teacher signals Portelas et al. (2020) Pateria et al. (2021); Hutsebaut-Buysse et al. (2022) or rewards Furelos-Blanco et al. (2023). These methods focus on the execution space, influencing reward signals for learning policies, but largely ignore the higher-order structures and relations between objects in the environment. Our work distinguishes itself by incorporating natural relational structures through knowledge graphs and leverages meta-level knowledge structures to guide interactions that govern low-level execution.
Meta-reasoning
Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, current meta-cognitive reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
Graph neural networks
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes. Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$. One typical task is to predict the label or value of the entire graph. Graph neural networks (GNN) (Gilmer et al. (2017); Hamilton et al. (2017) and others) use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$. Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
$h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable, permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which improves static attention problems in graph attention layers.
Policy gradient methods
We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992) , directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor critic methods using clipped objectives. We adapt PPO for learning policies.
3 Model
Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects. The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution. Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task generalization characteristics.
Figure 2: The meta-cognitive generalization framework consists of three levels: knowledge ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), which are based on abstraction and computation for providing generalization capabilities. The knowledge level contains type space knowledge graphs, which represent higher order abstractions. The instance space contains object instance graphs.The interaction level serves as an intermediary, making recommendations for interactions. The execution level is the lowest level at which the agent acts based on the recommended interaction subgoals. Computations proceed through recognition, recommendation, execution, and interactions to collect experiences for updates to interaction and meta policies.
Knowledge
The environment $\Xi$ is partially observable and consists of objects $e=\{e_{1},e_{2},...,e_{n}\}$ and an agent $A$. It is assumed that the agent has a recognition system that can access its partial view $X$ and associate objects with their corresponding types, such that the type of an object $e_{i}\in e$ is in entity types $E=\{E_{1},E_{2},...,E_{m}\}$. The agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ within the environment. When making decisions, the agent organizes its perceived knowledge into knowledge graphs: an instance-level graph and a type-level graph. An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. A type-level graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their type level or abstract relations $\Lambda$. Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another. The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and deriving their relationships. The function $\Psi:G_{I}\to G_{K}$ maps instance space graphs to type space graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects. If the agent encounters a novel object without a corresponding type, it must create a new type in its type space or assign the object to a related type. In the framework, $\Psi$ is learnable and represents a belief in mapping objects to types. However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
The agent’s type space level is also referred to as the meta-level. At the meta-level, the agent uses a meta-policy that accepts a type space knowledge graph and returns meta-level actions or interactions for object interaction. Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$. Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment. Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space knowledge graph and available interactions, while the meta-level knowledge $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space knowledge graph and interactions.
Interaction
After the agent constructs the knowledge graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s knowledge graphs and meta-level policy $\pi$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$. For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction. For $G_{K}$, the agent’s meta-policy returns the state value and interaction probability. The state value represents the value of activating that specific object. Then, an object $e_{i}$ for interaction is proposed through sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object. Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$. Acting on the recommended interaction goal, the agent begins interacting with the environment. After each round of interaction, the interaction status is passed to the meta-level, which then proposes the next candidate object and interaction. This process continues until the task is completed or the environment is terminated.
eXecution
Each agent has a single meta-level policy, whereas interaction policies are a collection of policies corresponding to different interactions. By following the interaction goals $g^{i}_{p}$ as directed by the meta-level, the agent interacts with objects. During an interaction $\lambda_{p}$, the agent executes a series of low-level actions $u$ on execution level states $s$ based on the relevant interaction policy $\mu_{p}$. For example, if there is a door, the agent can activate it and interact with it, like ‘open’ the door, even if this requires multiple low-level actions such as moving and toggling. As the agent interacts with objects, it collects trajectories of experiences at both the meta and interaction levels. The meta-level trajectory is used to learn the meta-level policy, while the low-level interaction experiences are used to learn the collection of interaction policies. The learning approach for the meta-policy and interaction policies is detailed below.
Meta-policy learning
Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge type encoding for different types of relationships. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,
$h_{v}^{\prime}=\alpha_{v,v}\theta h_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{u,v}\theta h_{u}$
where $h_{v}^{\prime}$ is the updated embedded representation of node $v$, $\mathcal{N}(v)$ is the set of neighbors of node $v$, and $h_{u}$ is the representation of the neighboring node. $\theta$ are learnable weights shared by source and target nodes in the neighborhood, and the attention coefficients $\alpha_{u,v}$ are computed using weighted node features.
The hidden representation after the final graph attention layer is pooled for graph rollout, which is then passed to the fully connected actor and critic layers of the model. The meta policy $\pi_{\theta}(.|G_{K})$, parameterized by $\theta$, returns a state value and an interaction or meta-action, $\lambda$, distribution. The recommender selects a target object and an interaction $\lambda_{p}$, and the agent works through low-level interactions based on guidance. At the end of the interaction, the status of the interaction is returned to the meta level, and the agent collects the trajectory of meta-level experiences, $\mathcal{D}_{\tau_{m}}$. An estimate of advantages is computed as $A_{m}=r_{m}+\gamma_{m}V_{\pi_{\theta}}(G_{K}^{\prime})-V_{\pi_{\theta}}(G_{K})$.
Meta level PPO optimizes a surrogate loss:
$\mathcal{L}_{m}^{\text{actor}}(\theta)=-\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\min\left(r_{\theta}A_{m},\text{clip}(r_{\theta},1-\epsilon,1+\epsilon)A_{m}\right)\right]$
with $r_{\theta}=\frac{\pi_{\theta}(.|G_{K})}{\pi_{\theta}^{r}(.|G_{K})}$ being the ratio of probabilities of actions at meta level roll-out. The critic loss is the mean squared error between the estimated values and the returns:
$\mathcal{L}_{m}^{\text{critic}}(\theta)=\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\frac{1}{2}(V_{\pi_{\theta}}(G_{K})-R_{m})^{2}\right]$
Considering the mean entropy $\mathcal{H}_{\theta}$, the overall minimization objective at the meta-level becomes:
$\mathcal{L}_{m}(\theta)=\mathcal{L}_{m}^{\text{actor}}(\theta)-\zeta_{m}^{h}\mathcal{H}_{\theta}+\zeta_{m}^{c}\mathcal{L}_{m}^{\text{critic}}(\theta)$
where $\zeta_{m}^{h}$, $\zeta_{m}^{c}$ are the entropy and value coefficients, respectively. The model parameters $\theta$ are optimized as $\theta\leftarrow\theta+\alpha_{m}\nabla_{\theta}L_{m}(\theta)$, and the updated meta-policy is used for the next episode of learning.
Interaction policy learning
Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately. When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$. The agent observes a low-level state $s$ from a convolutional neural network embedding, samples an action $u\sim\mu_{\phi_{p}}(.|s)$ using the respective interaction policy, acts in the environment, and collects the trajectory of experiences $\mathcal{D}_{\tau_{p}}$. The experiences collected are used to learn interaction policies. An estimate of advantages is computed as $A_{p}=r_{p}+\gamma_{p}V_{\mu_{\phi_{p}}}(s^{\prime})-V_{\mu_{\phi_{p}}}(s)$.
Each interaction level PPO optimizes a surrogate loss:
$\mathcal{L}_{p}^{\text{actor}}(\phi_{p})=-\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\min\left(r_{\phi_{p}}A_{p},\text{clip}(r_{\phi_{p}},1-\epsilon,1+\epsilon)A_{p}\right)\right]$
with $r_{\phi_{p}}=\frac{\mu_{\phi_{p}}(.|s)}{\mu_{\phi_{p}}^{r}(.|s)}$ is the ratio of probabilities of actions of policy and roll-out policy. The critic loss is the mean squared error between the estimated values and the returns:
$\mathcal{L}_{p}^{\text{critic}}(\phi_{p})=\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\frac{1}{2}(V_{\mu_{\phi_{p}}}(s)-R_{p})^{2}\right]$
Considering the mean entropy $\mathcal{H}_{\phi_{p}}$, the overall minimization objective becomes:
$\mathcal{L}_{p}(\phi_{p})=\mathcal{L}_{p}^{\text{actor}}(\phi_{p})-\zeta_{p}^{h}\mathcal{H}_{\phi_{p}}+\zeta_{p}^{c}\mathcal{L}_{p}^{\text{critic}}(\phi_{p})$
for each interaction $p$, where $\zeta_{p}^{h}$, $\zeta_{p}^{c}$ are the entropy and value coefficients, respectively. For each interaction policy, the model parameters $\phi_{p}$ are optimized as $\phi_{p}\leftarrow\phi_{p}+\alpha\nabla_{\phi_{p}}L_{p}(\phi_{p})$, and the updated interaction policies are used for the next episode of learning.
In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals. Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by meta-level knowledge guided object interactions, providing a mechanism for generalist behaviors.
4 Results and Discussion
Experiments
Our goal is to examine whether meta-level knowledge-guided interactions benefit agents in task generalization. To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks. MiniGrid world environments Chevalier-Boisvert et al. (2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes.
ObstructedMaze-Full
ObstructedMaze-1Dlhb
Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s performance on the designed tasks, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it. In the MiniGrid-ObstructedMaze-Full environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling. The default task is referred to as Task 0. Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
  • No context in training (Task 1) - a relation that the agent did not come across during training. It is simulated by placing the goal inside an empty box. During training, the blue ball was never in a box, although the agent encountered boxes and balls during training, but boxes contained keys.
  • Different context in testing (Task 2) - using a relation that the agent encountered during training but in a new context during testing. It is simulated by locking doors in the central room and placing keys directly in the room. During training, doors in the central room were never locked, nor was there direct access to keys.
  • Dynamic changes in the environment (Task 3) - when some event occurs, the environment changes. During training, the goal location remained unchanged. To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms. It is only done once, not repeatedly.
Agents learning to interact using type space knowledge
initialize model weights $\theta,\phi$ , parameters, $T$
for  $t=1$ to $T$ do
       /* collect exp for meta exp size K */
       for $k=1$ to $K$ do
             observe a partial view $X$
             make instance graph $G_{I}\leftarrow\Phi(X)$
             map $G_{I}$ to $G_{K}$, $G_{K}\leftarrow\Psi(G_{I})$
             recommend $(e_{i},\lambda_{p})\leftarrow\Upsilon(G_{K},G_{I},\pi)$
             interact with object $e_{i}$ following $\mu_{\lambda_{p}}$
             collect interaction exps $\mathcal{D}_{\tau_{p}}$ and compute $A_{p}$
       end for
      collect meta exp $\mathcal{D}_{\tau_{m}}$ and compute $A_{m}$
       /* Update policies */
       update $\pi$ by optimizing $\mathcal{L}_{m}(\theta)$
       for  $\lambda_{p}$ do
             update $\mu_{\lambda_{p}}$ by optimizing $\mathcal{L}_{p}(\phi_{p})$
       end for
end for
Algorithm 1 KIX procedure
We devised two variants: KIX-A and KIX-R. Both variants learn meta-policy and a collection of interaction policies. While the KIX-A variant combines both reaching objects and interacting with them, the KIX-R variant learns an additional interaction policy referred to as the ‘reachability’ policy, which is used to reach objects prior to interacting with them. All policies, including meta and interactions, are implemented as actor-critic PPO policies. Algorithm 1 provides an overview of the procedure. The agent’s partially observed view $X$ contains objects with identifiers and attributes like color and type. The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’. The ‘adjacent’ relation is derived using distance between the two objects. When the agent chooses an object to interact with, an ‘activated’ relation is added between the agent and the object node. The function $\Psi$ then constructs a type space graph from the instance space graph, encoding node types as attribute vectors. Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate both the meta-level and interactions such as ‘open’, ‘open with key’, ‘reveal’, ‘pickup’ (not shown for KIX-R), ‘drop’, and ‘reach’ in the ObstructedMaze-Full environment.
Agents learn how to perform low-level actions in the execution space using interaction policies. A convolutional network with three Conv2D layers with 16, 32, and 64 output channels, a max pooling layer added after the first convolution layer, and exponential linear units as activation functions is used to provide the input state embeddings for the interaction actor-critic. An activation indicator for the proposed interaction object is added as a new input channel, resulting in a four-channel input to the filters. Actor-critic networks are fully connected with a single hidden layer. Agents learn a meta-policy in the type space. Two graph attention layers with 4 heads and 16 dimensional embedding provide 64 dimensional input embeddings for the meta policy. During message passing, these layers share weights between source and target. The final graph attention layer is then used for graph readout, applying a global add pool. The meta-level actor and critic are both fully connected networks, with the critic learning a single value and the actor learning the likelihood of taking the meta-level actions ‘pickup’, ‘drop’, ‘reveal’, ‘open’, or ‘open with key’, each of which corresponds to an interaction policy.
Table 1: Training hyper-parameters
Parameter Value Parameter Value
meta exp size 8 gae lambda 0.95
exp size 64 entropy coeff 1e-04
meta minibatch size 64 value coeff 1.0
minibatch size 64 meta entropy coeff 1e-04
meta learning rate 1e-05 meta value coeff 1.0
learning rate 1e-05 ppo clip 0.2
gamma 0.99 ppo steps 4
The environment rewards the agents with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps. For interaction policies, the agent receives a reward of 1 when it has a successful interaction. At the meta-level, it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal. The base agent, BASE, has a similar actor-critic architecture and receives environment rewards. The training hyper-parameters are given in Table 1. Learning rates were determined using grid search with values 1e-3, 1e-4, 1e-5, and 1e-6. All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers.
As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and knowledge graphs, as indicated by higher meta-level returns. This, in turn, leads to better interaction experiences and further improvements in interaction policies. Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant. It suggests that the agents learn higher-level relations and apply these learned relations and interactions with objects to solve problems.
Meta-level knowledge-guided interactions aid agents
Our approach is motivated by the idea that meta-level guided interactions with objects help agents in generalizing while performing tasks, rather than acting directly at the execution level. Since the agent’s environmental return is determined by the number of steps taken to complete a task, completing tasks in fewer steps yields a higher return. Therefore, the distribution of environmental returns is a reliable indicator of the value of meta-level interaction recommendations for agents. We evaluate the distribution of environmental returns for each agent across four tasks. For each evaluation task, trained agents are run for 12K episodes by parallel workers, and their environmental rewards are tracked.
Figure 5: Evaluation return profiles of agents performing various tasks show that the proposed agents benefit from meta-level guided interactions.
To understand the effects of meta-guided interactions and ensure a fair comparison between agents, regardless of success rate differences, we examine the top $k$ returns with $k=100$. Figure 5 shows the return profiles of agents. It is evident from the distribution of returns that the proposed agents achieve higher returns on all tasks compared to the base agent, including dynamic tasks like Task 3. This demonstrates that meta-guided interactions not only assist agents in completing tasks in fewer steps, but also enable them to perform a wider range of tasks in both static and dynamic scenarios. Furthermore, higher returns of the second variant show that agents benefit from treating reachability as a separate interaction. In the absence of separate reachability, reaching is embedded into interactions. Separate reachability allows shared learning followed by interactions, which is particularly beneficial in dynamically shifting scenarios.
Transferability of policies
(a) Meta training
(b) Evaluation task returns
Figure 6: Evaluation return profiles show that interaction policies can be transferred from one environment to another.
The ObstructedMaze-1Dlhb environment is used to assess the transferability of interaction policies learned in one environment to another. In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze. The rooms are separated by a door, which is obstructed by a ball and locked. The keys are hidden inside boxes. Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies. During evaluation, agents use the learned meta-policy, but the interaction policies come from the ObstructedMaze-Full environment itself. Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants. It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
5 Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, type space knowledge graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects. Second, rather than directly acting in the environment, interactions with objects are critical in performing the tasks. To accomplish this, introduce an intermediary interaction level. The agent’s meta-level learns from type space knowledge graphs and makes recommendations on which objects to interact with and how to interact with them. In this sense, our model functions as an interaction goal-generating machine. Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries. We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization. Lastly, Our framework is modular and adaptable, allowing for the integration of knowledge structures and control variables at both the meta and interaction levels. This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.