### v3

# KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization

Abstract
People aptly exhibit general intelligence behaviors through flexible problem-solving and the ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time. In contrast, artificial agents tend to be specialists, lacking such generalist behaviors. To bridge this gap, artificial agents will require understanding and exploiting critical structured knowledge representations. We introduce a metacognitive reasoning framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects, by leveraging a type space, facilitate the learning of transferable interaction concepts and promote generalization. This framework offers a principled approach for integrating knowledge into reinforcement learning and holds promise as an enabler for generalist behaviors in artificial intelligence, robotics, and autonomous systems.
Introduction
The ability to actively reason, flexibly evaluate options and learn from experience is a hallmark of human intelligence. Core to this intelligence is the human ability to identify natural decompositions within complex environments and tasks with ease, enabling flexible and generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue necessitate a knowledge and interaction-centric metacognitive approach.
Meta-reasoning about next actions, their values, and possible next states underpins flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition. This form of reasoning extends beyond mere memory lookup. More broadly, meta-reasoning functions as a sense-making process within a knowledge system, enabling people to intuitively direct their learning to discern relations between entities, interact with the same entities in diverse ways, and ground these interactions for both short- and long-term goals. Metacognitive reasoning thus emerges as a foundational mechanism for generalization. In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems. While resource management is essential for efficient functioning, it is crucial to ideate metacognitive generalization, illustrated in Figure 1, to enable generalist behaviors in artificial agents.
(a) Overview
(b) Approach
Figure 1: a) Illustration of metacognitive reasoning process in a navigation domain, where an agent (depicted as a red triangle) must interpret its environment to achieve the goal of retrieving a blue ball located behind a locked door and various obstacles. This process requires reasoning over abstract object types and their relations, which in turn guide the agent’s interactions with objects. Such abstract reasoning fundamentally relies on higher-order abstract structures and compositional interactions over these abstractions. b) Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach exploits inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
Metacognitive reasoning is fundamentally grounded in interaction, requiring agents to draw upon their interpretations of abstract knowledge to guide their interactions. Interpretation precedes and shapes interaction; it involves developing both natural and abstract understandings of the environment that can be deployed as needed to make sense of the current state and plan subsequent interactions. To illustrate this process, consider the example in Figure 1a, which depicts an agent and its interactions in a navigation domain. To build knowledge of the environment, an agent must develop a conceptual understanding of the objects in the environment and how to interact with them. As some objects may even appear similar, the agent must decipher their abstract types. Learning to interact then involves learning which interactions are possible with these abstract object types and the relations between them. Type serves as a natural abstraction, and interpretation thus incorporates this meta-level reasoning over abstract knowledge. The utility of interactions is then learned through meta-level policies in this abstract space, which can subsequently be leveraged to recommend interactions and execute compositions of interactions at a lower level, thereby facilitating problem-solving.
Our proposal is based on three main principles for flexibly solving complex tasks. Natural abstraction: Real-world environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes. Entities in the instance space inherit properties and relations from a higher-level type space. Type-space graphs represent the underlying objects and their relations in environments. This abstraction is where the knowledge resides and it is the knowledge or meta level. Interactions: An agent’s metacognitive process employs the type-space graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals. While the metacognitive policy guides agents based on abstract knowledge, agents interact with entities in the instance space which are instantiations of abstract types. This is the interaction level. Composition of interactions: The agent interacts with meta-guided objects as a recommended interaction type and executes low-level policies corresponding to a guided interaction type. Therefore, a composition of these interactions as interaction pathways for graph traversal at both the meta level and the instance level facilitates execution of low-level actions which are loosely coupled with abstract knowledge through interactions.
Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents. Agents’ flexible behaviors stem from their ability to interpret knowledge, learn how to interact with objects through meta level policies and leverage interaction pathways enabling quick adaptation. The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains. We describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation in a navigation domain and discuss how we envision it enabling flexibility through meta-level guided interactions.
Background
In this section, we briefly describe closely related work and the general setting of meta-reasoning, graph neural networks, and policy gradient methods.
Related work
Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects. An object-centric model Veerapaneni et al. (2020) was presented in which object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning. Similarly, a symbolic object-centric representation Vijay et al. (2019) was used to study the impact of adding prior knowledge through relations in an object graph. These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making. Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks leveraging natural relational structures. We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is achieved through learning over knowledge graphs with neighborhood aggregation.
Hierarchical reinforcement learning approaches Dayan and Hinton (1992); Sutton et al. (1999); Pateria et al. (2021); Bagaria et al. (2021) learn a higher-level policy by choosing subtasks as higher-level actions, which decompose tasks into subtasks. Higher-level actions persist over a longer timescale, referred to as temporal abstraction Barto and Mahadevan (2003); Dietterich (2000); Sutton et al. (1999). In contrast to options Sutton et al. (1999), which combine a subtask with the core MDP, MAXQ value function decomposition Dietterich (2000) decomposes the core MDP into smaller sub-MDPs, each with a separate learnable policy. Other methods induce learning hierarchies based on teacher signals Portelas et al. (2020); Hutsebaut-Buysse et al. (2022) or a progression from easier to more difficult tasks Campero et al. (2020). These methods enable agents to learn hierarchy of policies, but they remain in the object space for both low- and high-level actions and do not meet a key meta-reasoning demand: type abstractions and loose coupling between abstraction and execution space. Methods such as Ten et al. (2022); Oudeyer et al. (2007) encourage exploration through intrinsic reward or motivation, Klyubin et al. (2005) develop internal goals for empowerment, focusing on aspects different than ours. On the other hand, neuro-symbolic reasoning approaches Hitzler and Sarker (2022); Zhang and Sheng (2024) aim to overcome the limitations of purely symbolic or purely neural methods by integrating them into hybrid models that combine the learning capacity of neural networks with the logical inference and interpretability of symbolic systems. However, significant challenges persist in facilitating interaction and integration between these disparate paradigms to ensure a seamless, flexible, and semantically grounded natural unification.
Our approach meets these demands by leveraging natural higher-order knowledge structures in the form of type space graphs for meta-level policy and facilitating interactions with objects via an intermediary between meta-level and execution space, thus enabling loose coupling and reuse of low-level policies. The appeal of our approach is that interaction policies can be swapped in and out depending on the interaction type. Our work distinguishes itself by incorporating relational abstract knowledge structures to guide interactions that govern low-level execution.
Meta-reasoning
Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, current meta-reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognitive approaches such as the one proposed here.
Graph neural networks
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is defined by a set of nodes $\mathcal{V}$ with node attributes $\mathtt{X}_{v}$ for $v\in\mathcal{V}$ and a set of edges $\mathcal{E}$ between nodes. Edge from node $u\in\mathcal{V}$ to node $v\in\mathcal{V}$ is given by $(u,v)\in\mathcal{E}$ with attributes $\mathtt{e}_{u,v}$. One typical task is to predict the label or value of the entire graph. Graph neural networks (GNN) (Gilmer et al. (2017); Hamilton et al. (2017) and others) use node connectivity as well as node and edge features to learn a permutation-invariant embedding vector for each node $h_{v},v\in\mathcal{V}$ and the entire graph $h_{\mathcal{G}}$. Typically, the representation is learned through message passing or neighborhood feature aggregation, in which node embedding vectors are updated in each layer based on message aggregation from neighbors:
$h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$ (1)
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which improves static attention problems in graph attention layers.
Policy gradient methods
We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches the state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992), directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor-critic methods using clipped objectives. We adapt PPO for learning policies.
Metacognitive framework
Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interpretation of knowledge and interpretation guided interactions with objects. The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution. Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task-generalization characteristics.
Figure 2: The meta-cognitive generalization framework consists of three levels: knowledge ($\mathcal{K}$), interaction ($\mathcal{I}$), and execution ($\mathcal{X}$), which are based on abstraction and computation for providing generalization capabilities. The knowledge level contains type space knowledge graphs, which represent higher order abstractions. The instance space contains object instance graphs. The interaction level serves as an intermediary, making recommendations for interactions. The execution level is the lowest level at which the agent acts based on the recommended interaction subgoals. Computations proceed through recognition, recommendation, execution, and interactions to collect experiences for updates to interaction and meta policies.
Knowledge
The environment $\Xi$ is partially observable and consists of objects $e=\{e_{1},e_{2},...,e_{n}\}$ and an agent $A$. It is assumed that the agent has a recognition system that can access its partial view $X$ and associate objects with their corresponding types, such that the type of an object $e_{i}\in e$ is in entity types $E=\{E_{1},E_{2},...,E_{m}\}$. The agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ within the environment. When making decisions, the agent organizes its perceived knowledge into graphs: an instance-level graph and a type-level graph. An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. A type-level graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their type level or abstract relations $\Lambda$. Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another. The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and deriving their relationships. The function $\Psi:G_{I}\to G_{K}$ maps instance space graphs to type space graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects. If the agent encounters a novel object without a corresponding type, it must create a new type in its type space or assign the object to a related type. In the framework, $\Psi$ is learnable and represents a belief in mapping objects to types. However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
The agent’s knowledge level is also referred to as the meta-level. At the meta-level, the agent uses a meta-policy that accepts a type space graph and returns interactions or meta-level actions for object interaction. Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$. Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment. Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space graph and available interactions, while the meta or knowledge level $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space graph and meta-level actions or interactions.
Interaction
After the agent constructs the type and instance graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s type and instance graphs and meta-level policy $\pi$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$. For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction. For $G_{K}$, the agent’s meta-policy returns the state value and the interaction probability. The state value represents the value of activating that specific object. Then, an object $e_{i}$ for interaction is proposed by sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object. Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$. Acting on the recommended interaction goal, the agent begins interacting with the environment. After each round of interaction, the interaction status is passed to the meta-level, which then proposes the next candidate object and interaction. This process continues until the task is completed or the environment is terminated.
eXecution
Each agent has a meta-level policy, whereas interaction policies are a collection of policies corresponding to different interactions. By following the interaction goals $g^{i}_{p}$ as directed by the meta-level, the agent interacts with objects. During an interaction $\lambda_{p}$, the agent executes a series of low-level actions $u$ on execution level states $s$ based on the relevant interaction policy $\mu_{p}$. For example, if there is a door, the agent can activate it and interact with it, like ‘open’ the door, even if this requires multiple low-level actions such as moving and toggling. As the agent interacts with objects, it collects trajectories of experiences at both the meta and interaction levels. The meta-level trajectory is used to learn the meta-level policy, while the low-level interaction experiences are used to learn the collection of interaction policies. The learning approach for the meta-policy and interaction policies is detailed below.
Meta-policy learning
Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge-type encoding for different types of relations. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,
$h_{v}^{\prime}=\alpha_{v,v}\theta h_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{u,v}\theta h_{u}$ (2)
where $h_{v}^{\prime}$ is the updated embedded representation of node $v$, $\mathcal{N}(v)$ is the set of neighbors of node $v$, and $h_{u}$ is the representation of the neighboring node. $\theta$ are learnable weights shared by source and target nodes in the neighborhood, and the attention coefficients $\alpha_{u,v}$ are computed using weighted node features.
The hidden representation after the final graph attention layer is pooled for graph rollout, which is then passed to the fully connected actor and critic layers of the model. The meta policy $\pi_{\theta}(.|G_{K})$, parameterized by $\theta$, returns a state value and an interaction or meta-action, $\lambda$, distribution. The recommender selects a target object and an interaction $\lambda_{p}$, and the agent works through low-level interactions based on guidance. At the end of the interaction, the status of the interaction is returned to the meta level, and the agent collects the trajectory of meta-level experiences, $\mathcal{D}_{\tau_{m}}$. An estimate of advantages is computed as $A_{m}=r_{m}+\gamma_{m}V_{\pi_{\theta}}(G_{K}^{\prime})-V_{\pi_{\theta}}(G_{K})$.
Meta level PPO optimizes a surrogate loss:
$\mathcal{L}_{m}^{\text{actor}}(\theta)=-\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\min\left(r_{\theta}A_{m},\text{clip}(r_{\theta},1-\epsilon,1+\epsilon)A_{m}\right)\right]$ (3)
with $r_{\theta}=\frac{\pi_{\theta}(.|G_{K})}{\pi_{\theta}^{r}(.|G_{K})}$ being the ratio of probabilities of actions at meta level roll-out. The critic loss is the mean squared error between the estimated values and the returns:
$\mathcal{L}_{m}^{\text{critic}}(\theta)=\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\frac{1}{2}(V_{\pi_{\theta}}(G_{K})-R_{m})^{2}\right]$ (4)
Considering the mean entropy $\mathcal{H}_{\theta}$, the overall minimization objective at the meta-level becomes:
$\mathcal{L}_{m}(\theta)=\mathcal{L}_{m}^{\text{actor}}(\theta)-\zeta_{m}^{h}\mathcal{H}_{\theta}+\zeta_{m}^{c}\mathcal{L}_{m}^{\text{critic}}(\theta)$ (5)
where $\zeta_{m}^{h}$, $\zeta_{m}^{c}$ are the entropy and value coefficients, respectively. The model parameters $\theta$ are optimized as $\theta\leftarrow\theta+\alpha_{m}\nabla_{\theta}L_{m}(\theta)$, and the updated meta-policy is used for the next episode of learning.
Interaction policy learning
Interaction policies are a repository of policies for different interactions which can be recommended by the meta level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately. When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$. The agent observes a low-level state $s$ from a convolutional neural network embedding, samples an action $u\sim\mu_{\phi_{p}}(.|s)$ using the respective interaction policy, acts in the environment, and collects the trajectory of experiences $\mathcal{D}_{\tau_{p}}$. The experiences collected are used to learn interaction policies. An estimate of advantages is computed as $A_{p}=r_{p}+\gamma_{p}V_{\mu_{\phi_{p}}}(s^{\prime})-V_{\mu_{\phi_{p}}}(s)$.
Each interaction level PPO optimizes a surrogate loss:
$\mathcal{L}_{p}^{\text{actor}}(\phi_{p})=-\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\min\left(r_{\phi_{p}}A_{p},\text{clip}(r_{\phi_{p}},1-\epsilon,1+\epsilon)A_{p}\right)\right]$ (6)
with $r_{\phi_{p}}=\frac{\mu_{\phi_{p}}(.|s)}{\mu_{\phi_{p}}^{r}(.|s)}$ is the ratio of probabilities of actions of policy and roll-out policy. The critic loss is the mean squared error between the estimated values and the returns:
$\mathcal{L}_{p}^{\text{critic}}(\phi_{p})=\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\frac{1}{2}(V_{\mu_{\phi_{p}}}(s)-R_{p})^{2}\right]$ (7)
Considering the mean entropy $\mathcal{H}_{\phi_{p}}$, the overall minimization objective becomes:
$\mathcal{L}_{p}(\phi_{p})=\mathcal{L}_{p}^{\text{actor}}(\phi_{p})-\zeta_{p}^{h}\mathcal{H}_{\phi_{p}}+\zeta_{p}^{c}\mathcal{L}_{p}^{\text{critic}}(\phi_{p})$ (8)
for each interaction $p$, where $\zeta_{p}^{h}$, $\zeta_{p}^{c}$ are the entropy and value coefficients, respectively. For each interaction policy, the model parameters $\phi_{p}$ are optimized as $\phi_{p}\leftarrow\phi_{p}+\alpha\nabla_{\phi_{p}}L_{p}(\phi_{p})$, and the updated interaction policies are used for the next episode of learning.
In summary, when viewing tasks as meta-policy guided interactions with objects, problems are naturally decomposed into interaction goals. Unlike the models that are tightly coupled with low-level representations through end-to-end policies, in the proposed framework, agent actions are aided by the interpretation of meta-level knowledge to guide object interactions, providing a mechanism for generalist behaviors.
Results and discussion
Experiments
Our goal is to examine whether meta-level knowledge-guided interactions benefit agents in task generalization. To assess the capabilities of the proposed agents, we need a challenging environment having ingredients for higher-order reasoning to solve tasks. MiniGrid world navigation environments Chevalier-Boisvert et al. (2018) are suitable for this purpose; they are procedurally generated environments where the agent interacts with objects such as doors, keys, balls, and boxes. The obstructed maze environments, shown in Figure 3, are particularly challenging due to partial observability, stochasticity, and sparse rewards, with only a few claims for solving it. In the MiniGrid-ObstructedMaze-Full environment, the agent is initially placed in the middle room, with a 7x7 view. At the low level, the agent can move forward, left, right, pick, drop, and toggle, and its goal is to pick up the blue ball, which is randomly placed in one of the corner rooms. Some of the doors are locked and require a key of the same color to open. Keys are hidden inside boxes, but the agent cannot see inside the boxes without toggling.
(a) ObstructedMaze-Full
(b) ObstructedMaze-1Dlhb
Figure 3: Environments: ObstructedMaze-Full is used to experimentally evaluate the agent’s performance on the designed tasks, while ObstructedMaze-1Dlhb is used to evaluate the transferability of interaction policies.
The default task is referred to as Task 0. Furthermore, we identify the following new task scenarios, which were presented to the agent during testing but not during training:
  • No context in training (Task 1) - a relation that the agent did not come across during training. It is simulated by placing the goal inside an empty box. During training, the blue ball was never in a box, although the agent encountered boxes and balls during training, but boxes contained keys.
  • Different context in testing (Task 2) - using a relation that the agent encountered during training but in a new context during testing. It is simulated by locking doors in the central room and placing keys directly in the room. During training, doors in the central room were never locked, nor was there direct access to keys.
  • Dynamic changes in the environment (Task 3) - when some event occurs, the environment changes. During training, the goal location remained unchanged. To simulate, as soon as the agent enters the room with the goal, relocate the goal to one of the neighboring rooms. It is only done once, not repeatedly.
Learning to interact using type space knowledge
We devised two variants: KIX-A and KIX-R. Both variants learn a meta-policy and a collection of interaction policies. While the KIX-A variant combines both reaching objects and interacting with them, the KIX-R variant learns an additional interaction policy referred to as the ‘reachability’ policy, which is used to reach objects prior to interacting with them. All policies, including meta and interactions, are implemented as actor-critic PPO policies.
initialize model weights $\theta,\phi$ , parameters, $T$
for  $t=1$ to $T$ do
   /* collect meta experiences */
 for $k=1$ to $K$ do
      observe a partial view $X$
      make instance graph $G_{I}\leftarrow\Phi(X)$
      map $G_{I}$ to $G_{K}$, $G_{K}\leftarrow\Psi(G_{I})$
      recommend $(e_{i},\lambda_{p})\leftarrow\Upsilon(G_{K},G_{I},\pi)$
      interact with object $e_{i}$ following $\mu_{\lambda_{p}}$
      collect interaction exps $\mathcal{D}_{\tau_{p}}$ and compute $A_{p}$
   end for
  collect meta exp $\mathcal{D}_{\tau_{m}}$ and compute $A_{m}$
   /* Update policies */
   update $\pi$ by optimizing $\mathcal{L}_{m}(\theta)$
 for  $\lambda_{p}$ do
      update $\mu_{\lambda_{p}}$ by optimizing $\mathcal{L}_{p}(\phi_{p})$
   end for
end for
Algorithm 1 KIX procedure
Algorithm 1 provides an overview of the procedure. The agent’s partially observed view $X$ contains objects with identifiers and attributes like color and type. The function $\Phi$ takes $X$ and creates an instance space graph by using visible objects as nodes, including the agent, and determining relationships between nodes such as ‘visible’, ‘adjacent’, and ‘carrying’. The ‘adjacent’ relation is derived using distance between the two objects. When the agent chooses an object to interact with, an ‘activated’ relation is added between the agent and the object node. The function $\Psi$ then constructs a type space graph from the instance space graph, encoding node types as attribute vectors. Edges in a type space graph have the same meaning as those in an instance space graph and are encoded as one-hot encoded edge attribute vectors.
Agents learn how to perform low-level actions in the execution space using interaction policies. A convolutional network with three Conv2D layers with 16, 32, and 64 output channels, a max pooling layer added after the first convolution layer, and exponential linear units as activation functions is used to provide the input state embeddings for the interaction actor-critic. An activation indicator for the proposed interaction object is added as a new input channel, resulting in a four-channel input to the filters. Actor-critic networks are fully connected with a single hidden layer. Agents learn a meta-policy in the type space. Two graph attention layers with 4 heads and 16 dimensional embedding provide 64 dimensional input embeddings for the meta policy. During message passing, these layers share weights between source and target. The final graph attention layer is then used for graph readout, applying a global add pool. The meta-level actor and critic are both fully connected networks, with the critic learning a single value and the actor learning the likelihood of taking the meta-level actions ‘pickup’, ‘drop’, ‘reveal’, ‘open’, or ‘open with key’, each of which corresponds to an interaction policy.
Table 1: Training hyper-parameters
Parameter Value Parameter Value
meta exp size 8 gae lambda 0.95
exp size 64 entropy coeff 1e-04
meta minibatch size 64 value coeff 1.0
minibatch size 64 meta entropy coeff 1e-04
meta learning rate 1e-05 meta value coeff 1.0
learning rate 1e-05 ppo clip 0.2
gamma 0.99 ppo steps 4
The environment rewards the agents with $r=1-(c/c_{T})$ where $c$ is the number of steps taken and $c_{T}$ is the max possible steps. For interaction policies, the agent receives a reward of 1 when it has a successful interaction. At the meta-level, it is rewarded with 0.1 for a successful proposed interaction and $1+r$ for achieving the environment goal. The base agent, BASE, is an end to end policy learner with a similar actor-critic architecture and receives environment rewards. The training hyper-parameters are given in Table 1. Learning rates were determined using grid search with values 1e-3, 1e-4, 1e-5, and 1e-6. All the agents are trained for $\sim 6.2\times 10^{8}$ steps, using 32 parallel workers in a high performance computing cluster.
Figure 4: Agents learning to interact: The learning progress of the agents KIX-A (blue) and KIX-R (red) is shown as returns over steps, which illustrate learning both the meta-level and interactions such as ‘reach’ (only applicable for KIX-R), ‘open’, ‘open with key’, ‘reveal’, ‘pickup’, ‘drop’ (not shown) in the ObstructedMaze-Full environment.
As shown in Figure 4, as training progresses, agents learn to interact with objects and associate higher-level structured knowledge with these interactions using object types and type and instance space graphs, as indicated by higher meta-level returns. This, in turn, leads to better interaction experiences and further improvements in interaction policies. Furthermore, the KIX-R variant first focuses on learning how to reach different objects, but once it learns to reach, it begins to learn interactions in fewer steps and with much higher returns compared to the KIX-A variant. It suggests that the agents learn higher-order relations and apply these learned relations and interactions with objects to solve problems.
Meta-level guided interactions aid agents
Our approach is motivated by the idea that meta-level guided interactions with objects help agents in generalizing while performing tasks, rather than acting directly at the execution level. Since the agent’s environmental return is determined by the number of steps taken to complete a task, completing tasks in fewer steps yields a higher return. Therefore, the distribution of environmental returns is a reliable indicator of the value of meta-level interaction recommendations for agents. We evaluate the distribution of environmental returns for each agent across four tasks. For each evaluation task, trained agents are run for 12K episodes by parallel workers, and their environmental rewards are tracked.
Figure 5: Evaluation return profiles of agents performing various tasks show that the proposed agents benefit from meta-level guided interactions.
To understand the effects of meta-guided interactions and ensure a fair comparison between agents, regardless of success rate differences, we examine the top $k$ returns with $k=100$. Figure 5 shows the return profiles of agents. It is evident from the distribution of returns that the proposed agents achieve higher returns on all tasks compared to the base agent, including dynamic tasks like Task 3. This demonstrates that meta-guided interactions not only assist agents in completing tasks in fewer steps, but also enable them to perform a wider range of tasks in both static and dynamic scenarios. Furthermore, higher returns of the second variant show that agents benefit from treating reachability as a separate interaction. In the absence of separate reachability, reaching is embedded into interactions. Separate reachability allows shared learning followed by interactions, which is particularly beneficial in dynamically shifting scenarios.
Transferability of policies
(a) Meta training
(b) Evaluation of task returns
Figure 6: Evaluation return profiles show that interaction policies can be transferred from one environment to another.
The ObstructedMaze-1Dlhb environment is used to assess the transferability of interaction policies learned in one environment to another. In this environment, the agent’s goal is to pick up a blue ball hidden in a two-room maze. The rooms are separated by a door that is obstructed by a ball and locked. The keys are hidden inside boxes. Agents employ previously acquired interaction policies from the base task in the ObstructedMaze-Full environment while only learning the meta-policies. During evaluation, agents use the learned meta-policy, but the interaction policies are from the ObstructedMaze-Full environment itself. Figure 6 shows the meta training returns and evaluation return profiles, similar to the main environment, for both agent variants. It is evident that agents have gained the ability to exploit interaction policies from the previous environment, and the second variant, which employs a separate reachability policy, appears to have benefited more.
Conclusion
In this work, we proposed a metacognitive framework that promotes generalist behaviors through three main contributions. First, generalist behaviors necessitate abstraction and composition; however, current models lack natural abstractions and compositions that can preserve both object abstractions and relational knowledge structures. In our model, type-space graphs are used to represent and ground natural abstractions and compositions while also preserving relational knowledge between objects. Second, rather than acting directly in the environment, interactions with objects are critical in performing tasks. To accomplish this, we introduced an intermediary interaction level. The agent’s meta-level interprets abstract relational knowledge, learns a meta-policy using type-space graphs, and makes recommendations on which objects to interact with and how to interact with them. In this sense, our model functions as an interpretation-based interaction goal-generating machine. Third, interactions employing interaction policies enable the necessary loose coupling required for generalization by acting as intermediaries. We demonstrated that metacognitive knowledge-guided interactions aid agents in generalization. Lastly, Our framework is modular and adaptable, enabling the integration of knowledge structures and control variables at both the meta and interaction levels. This framework is promising, particularly for domains where higher-order knowledge and interaction compositions can be exploited, such as robotics, autonomous and cognitive systems.